In a triangle ABC if s-a/11=s-b/12=s-c/13 then Tan^2(A/2) is equal to.

Grade:11

1 Answers

Jayanth
6 years ago
Tan^2(A/2) = (s-b)(s-c)/s.(s-a).............eq 1Let s-a/11 = s-b/12 = s-c/13 = kSo, s-a = 11k, s-b = 12k, s-c = 13k(s-a) + (s-b) + (s-c) = 36ks + 2s - (a+b+c) = 36kS= 36k...Substitute these values in eq 1Tan^2(A/2) = 12k.13k/36k.11k = 13/33